# Generalization of DNNs:

### v1

@@ 3.2 Deep Networks
Since an AccNet is simply the composition of multiple shallow nets, the functions represented by an AccNet is included in the set of composition of $F_{1}$ balls. More precisely, if $\left\|W_{\ell}\right\|^{2}+\left\|V_{\ell}\right\|^{2}+\left\|b_{\ell}\right\|^{2}\leq 2R_{\ell}$ then $f_{L:1}$ belongs to the set $\left\{g_{L}\circ\cdots\circ g_{1}:g_{\ell}\in B_{F_{1}}(0,R_{\ell})\right\}$ for some $R_{\ell}$, which is width agnostic.
As already noticed in [7], the covering number number is well-behaved under composition, this allows us to bound the complexity of AccNets in terms of the individual shallow nets it is made up of:
@@
Consider an accordion net of depth $L$ and widths $d_{L},\dots,d_{0}$, with corresponding set of functions =$\mathcal{F}=\{f_{L:1}:\left\|f_{\ell}\right\|_{F_{1}}\leq R_{\ell},\text{Lip}(f_{\ell})\leq\rho_{\ell}\}$. With probability $1-\delta$ over the sampling of the training set $X$ from the distribution $\pi$ supported in $B(0,b)$, we have for all $f\in\mathcal{F}$
$\mathcal{L}(f)-\tilde{\mathcal{L}}_{N}(f)$ $\leq C\rho b\rho_{L:1}\sum_{\ell=1}^{L}\frac{R_{\ell}}{\rho_{\ell}}\sqrt{d_{\ell}+d_{\ell-1}}\frac{\log N}{\sqrt{N}}(1+o(1))+c_{0}\sqrt{\frac{2\log\nicefrac{{2}}{{\delta}}}{N}}.$
@@
The Lipschitz constants $Lip(f_{\ell})$ are difficult to compute exactly, so it is easiest to simply bound it by the product of the operator norms $Lip(f_{\ell})\leq\left\|W_{\ell}\right\|_{op}\left\|V_{\ell}\right\|_{op}$, but this bound can often be quite loose. The fact that our bound depends on the Lipschitz constants rather than the operator norms $\left\|W_{\ell}\right\|_{op},\left\|V_{\ell}\right\|_{op}$ is thus a significant advantage.
This bound can be applied to a FCNNs with weight matrices $M_{1},\dots,M_{L+1}$, by replacing the middle $M_{\ell}$ with SVD decomposition $USV^{T}$ in the middle by two matrices $W_{\ell-1}=\sqrt{S}V^{T}$ and $V_{\ell}=U\sqrt{S}$, so that the dimensions can be chosen as the rank $d_{\ell}=\mathrm{Rank}M_{\ell+1}$. The Frobenius norm of the new matrices equal the nuclear norm of the original one $\left\|W_{\ell-1}\right\|_{F}^{2}=\left\|V_{\ell}\right\|_{F}^{2}=\left\|M_{\ell}\right\|_{*}$. Some bounds assuming rank sparsity of the weight matrices also appear in [43]. And several recent results have shown that weight-decay leads to a low-rank bias on the weight matrices of the network [27, 26, 19] and replacing the Frobenius norm regularization with a nuclear norm regularization (according to the above mentioned equivalence) will only increase this low-rank bias.
We compute in Figure 1 the upper bound of Theorem 2 for both AccNets and DNNs, and even though we observe a very large gap (roughly of order $10^{3}$), we do observe that it captures rate/scaling of the test error (the log-log slope) well. So this generalization bound could be well adapted to predicting rates, which is what we will do in the next section.
@@
Figure 1: Visualization of scaling laws. We observe that deep networks (either AccNets or DNNs) achieve better scaling laws than kernel methods or shallow networks on certain compositional tasks, in agreement with our theory. We also see that our new generalization bounds approximately recover the right saling laws (even though they are orders of magnitude too large overall). We consider a compositional true function $f^{*}=h\circ g$ where $g$ maps from dimension 15 to 3 while h maps from 3 to 20, and we denote $\nu_{g},\nu_{h}$ for the number of times $g,h$ are differentiable. In the first plot $\nu_{g}=8,\nu_{h}=1$ so that $g$ is easy to learn while $h$ is hard, whereas in the second plot $\nu_{g}=9,\nu_{h}=9$, so both $g$ and $h$ are relatively easier. The third plot presents the decay in test error and generalization bounds for networks evaluated using the real-world dataset, WESAD [38].
@@ 4.1 Composition of Sobolev Balls
The family of Sobolev norms capture some notion of regularity of a function, as it measures the size of its derivatives. The Sobolev norm of a function $f:\mathbb{R}^{d_{in}}\to\mathbb{R}$ is defined in terms of its derivatives $\partial_{x}^{\alpha}f$ for some $d_{in}$-multi-index $\alpha$, namely the $W^{\nu,p}(\pi)$-Sobolev norm with integer $\nu$ and $p\geq 1$ is defined as
@@
Given a distribution $\pi$ with support the $L_{2}$ ball with radius b, we have that with probability $1-\delta$ for all functions $f\in\mathcal{F}=\left\{f:\left\|f\right\|_{W^{\nu,2}}\leq R,\left\|f\right\|_{\infty}\leq R\right\}$
$\mathcal{L}(f)-\tilde{\mathcal{L}}_{N}(f)\leq 2\rho C_{1}RE_{\nicefrac{{\nu}}{{d}}}(N)+c_{0}\sqrt{\frac{2\log\nicefrac{{2}}{{\delta}}}{N}}.$
where $E_{r}(N)=N^{-\frac{1}{2}}$ if $r>\frac{1}{2}$, $E_{r}(N)=N^{-\frac{1}{2}}\log N$ if $r=\frac{1}{2}$, and $E_{r}(N)=N^{-r}$ if $r<\frac{1}{2}$.
But this result also illustrates the curse of dimensionality: the differentiability $\nu$ needs to scale with the input dimension $d_{in}$ to obtain a reasonable rate. If instead $\nu$ is constant and $d_{in}$ grows, then the number of datapoints $N$ needed to guarantee a generalization gap of at most $\epsilon$ scales exponentially in $d_{in}$, i.e. $N\sim\epsilon^{-\frac{d_{in}}{\nu}}$. One way to interpret this issue is that regularity becomes less and less useful the larger the dimension: knowing that similar inputs have similar outputs is useless in high dimension where the closest training point $x_{i}$ to a test point $x$ is typically very far away.
Figure 2: A comparison of empirical and theoretical error rates. The first plot illustrates the log decay rate of the test error with respect to the dataset size $N$ based on our empirical simulations. The second plot depicts the theoretical decay rate of the test error as discussed in Section 4.1, $-\min\{\frac{1}{2},\frac{\nu_{g}}{d_{in}},\frac{\nu_{h}}{d_{mid}}\}$. The final plot on the right displays the difference between the two. The lower left region represents the area where $g$ is easier to learn than $h$, the upper right where $h$ is easier to learn than $g$, and the lower right region where both $f$ and $g$ are easy.
@@ 4.1.1 Breaking the Curse of Dimensionality with Compositionality
Manifold hypothesis: If the input distribution lies on a $d_{surf}$-dimensional manifold, the error rates typically depends on $d_{surf}$ instead of $d_{in}$ [39, 10].
Figure 3: Comparing error rates for shallow and AccNets: shallow nets vs. AccNets, and kernel methods vs. AccNets. The left two graphs shows the empirical decay rate of test error with respect to dataset size (N) for both shallow nets and kernel methods. In contrast to our earlier empirical findings for AccNets, both shallow nets and kernel methods exhibit a slower decay rate in test error. The right two graphs present the differences in log decay rates between shallow nets and AccNets, as well as between kernel methods and AccNets. AccNets almost always obtain better rates, with a particularly large advantage at the bottom and middle-left.
Known Symmetries: If $f^{*}(g\cdot x)=f^{*}(x)$ for a group action $\cdot$ w.r.t. a group $G$, then $f^{*}$ can be written as the composition of a modulo map $g^{*}:\mathbb{R}^{d_{in}}\to\nicefrac{{\mathbb{R}^{d_{in}}}}{{G}}$ which maps pairs of inputs which are equivalent up to symmetries to the same value (pairs $x,y$ s.t. $y=g\cdot x$ for some $g\in G$), and then a second function $h^{*}:\nicefrac{{\mathbb{R}^{d_{in}}}}{{G}}\to\mathbb{R}^{d_{out}}$, then the complexity of the task will depend on the dimension of the modulo space $\nicefrac{{\mathbb{R}^{d_{in}}}}{{G}}$ which can be much lower. If the symmetry is known, then one can for example fix $g^{*}$ and only learn $h^{*}$ (though other techniques exist, such as designing kernels or features that respect the same symmetries) [31].
Symmetry Learning: However if the symmetry is not known then both $g^{*}$ and $h^{*}$ have to be learned, and this is where we require feature learning and/or compositionality. Shallow networks are able to learn translation symmetries, since they can learn so-called low-index functions which satisfy $f^{*}(x)=f^{*}(Px)$ for some projection $P$ (with a statistical complexity that depends on the dimension of the space one projects into, not the full dimension [5, 2]). Low-index functions correspond exactly to the set of functions that are invariant under translation along the kernel $\ker P$. To learn general symmetries, one needs to learn both $h^{*}$ and the modulo map $g^{*}$ simultaneously, hence the importance of feature learning.
For $g^{*}$ to be learnable efficiently, it needs to be regular enough to not suffer from the curse of dimensionality, but many traditional symmetries actually have smooth modulo maps, for example the modulo map $g^{*}(x)=\left\|x\right\|^{2}$ for rotation invariance. This can be understood as a special case of composition of Sobolev functions, whose generalization gap can be bounded:
@@
Consider the function set $\mathcal{F}=\mathcal{F}_{L}\circ\cdots\circ\mathcal{F}_{1}$ where =$\mathcal{F}_{\ell}=\left\{f_{\ell}:\mathbb{R}^{d_{\ell-1}}\to\mathbb{R}^{d_{\ell}}\text{ s.t. }\left\|f_{\ell}\right\|_{W^{\nu_{\ell},2}}\leq R_{\ell},\left\|f_{\ell}\right\|_{\infty}\leq b_{\ell},Lip(f_{\ell})\leq\rho_{\ell}\right\}$, and let $r_{min}=\min_{\ell}r_{\ell}$ for $r_{\ell}=\frac{\nu_{\ell}}{d_{\ell-1}}$, then with probability $1-\delta$ we have for all $f\in\mathcal{F}$
$\mathcal{L}(f)-\tilde{\mathcal{L}}_{N}(f)$ $\leq\rho C_{0}\left(\sum_{\ell=1}^{L}\left(C_{\ell}\rho_{L:\ell+1}R_{\ell}\right)^{\frac{1}{r_{min}+1}}\right)^{r_{min}+1}E_{r_{min}}(N)+c_{0}\sqrt{\frac{2\log\nicefrac{{2}}{{\delta}}}{N}},$
where $C_{\ell}$ depends only on $d_{\ell-1},d_{\ell},\nu_{\ell},b_{\ell-1}$.
We see that only the smallest ratio $r_{min}$ matters when it comes to the rate of learning. And actually the above result could be slightly improved to show that the sum over all layers could be replaced by a sum over only the layers where the ratio $r_{\ell}$ leads to the worst rate $E_{r_{\ell}}(N)=E_{r_{min}}(N)$ (and the other layers contribute an asymptotically subdominant amount).
Coming back to the symmetry learning example, we see that the hardness of learning a function of the type $f^{*}=h\circ g$ with inner dimension $d_{mid}$ and regularities $\nu_{g}$ and $\nu_{h}$, the error rate will be (up to log terms) $N^{-\min\{\frac{1}{2},\frac{\nu_{g}}{d_{in}},\frac{\nu_{h}}{d_{mid}}\}}$. This suggests the existence of three regimes depending on which term attains the minimum: a regime where both $g$ and $h$ are easy to learn and we have $N^{-\frac{1}{2}}$ learning, a regime $g$ is hard, and a regime where $h$ is hard. The last two regimes differentiate between tasks where learning the symmetry is hard and those where learning the function knowing its symmetries is hard.
In contrast, without taking advantage of the compositional structure, we expect $f^{*}$ to be only $\min\{\nu_{g},\nu_{h}\}$ times differentiable, so trying to learn it as a single Sobolev function would lead to an error rate of $N^{-\min\{\frac{1}{2},\frac{\min\{\nu_{g},\nu_{h}\}}{d_{in}}\}}=N^{-\min\{\frac{1}{2},\frac{\nu_{g}}{d_{in}},\frac{\nu_{h}}{d_{in}}\}}$ which is no better than the compositional rate, and is strictly worse whenever $\nu_{h}<\nu_{g}$ and $\frac{\nu_{h}}{d_{in}}<\frac{1}{2}$ (we can always assume $d_{mid}\leq d_{in}$ since one could always choose $d=id$).
Furthermore, since multiple compositions are possible, one can imagine a hierarchy of symmetries that slowly reduce the dimensionality with less and less regular modulo maps. For example one could imagine a composition $f_{L}\circ\cdots\circ f_{1}$ with dimensions $d_{\ell}=d_{0}2^{-\ell}$ and regularities $\nu_{\ell}=d_{0}2^{-\ell}$ so that the ratios remain constant $r_{\ell}=\frac{d_{0}2^{-\ell}}{d_{0}2^{-\ell+1}}=\frac{1}{2}$, leading to an almost parametric rate of $N^{-\frac{1}{2}}\log N$ even though the function may only be $d_{0}2^{-L}$ times differentiable. Without compositionality, the rate would only be $N^{-2^{-L}}$.
@@
In the case of a single Sobolev function, one can show that the rate $E_{\nicefrac{{\nu}}{{d}}}(N)$ is in some sense optimal, by giving an information theoretic lower bound with matching rate. A naive argument suggests that the rate of $E_{\min\{r_{1},\dots,r_{L}\}}(N)$ should similarly be optimal: assume that the minimum $r_{\ell}$ is attained at a layer $\ell$, then one can consider the subset of functions such that the image $f_{\ell-1:1}(B(0,r))$ contains a ball $B(z,r^{\prime})\subset\mathbb{R}^{d_{\ell-1}}$ and that the function $f_{L:\ell+1}$ is $\beta$-non-contracting $\left\|f_{L:\ell+1}(x)-f_{L:\ell+1}(y)\right\|\geq\beta\left\|x-y\right\|$, then learning $f_{L:1}$ should be as hard as learning $f_{\ell}$ over the ball $B(z,r^{\prime})$ (more rigorously this could be argued from the fact that any $\epsilon$-covering of $f_{L:1}$ can be mapped to an $\nicefrac{{\epsilon}}{{\beta}}$-covering of $f_{\ell}$), thus forcing a rate of at least $E_{r_{\ell}}(N)=E_{\min\{r_{1},\dots,r_{L}\}}(N)$.
An analysis of minimax rates in a similar setting has been done in [22].
Remark.
The type of model that seems most appropriate to learning composition of Sobolev functions are arguably Deep Gaussian Processes (DGP). While it should be relatively straightforward to use DGP with a Lipschitzness constraint to recover optimal rates when the dimensions $d_{\ell}$ and differentiabilities $\nu_{\ell}$ of the true functions are known in advance, it might not possible in general. We will show that AccNet can recover almost optimal rates thanks to the adaptivity of the $F_{1}$ norm.

### v2

@@ 3.2 Deep Networks
Since an AccNet is simply the composition of multiple shallow nets, the functions represented by an AccNet is included in the set of composition of $F_{1}$ balls. More precisely, if $\left\|W_{\ell}\right\|^{2}+\left\|V_{\ell}\right\|^{2}+\left\|b_{\ell}\right\|^{2}\leq 2R_{\ell}$ then $f_{L:1}$ belongs to the set $\left\{g_{L}\circ\cdots\circ g_{1}:g_{\ell}\in B_{F_{1}}(0,R_{\ell})\right\}$ for some $R_{\ell}$.
As noticed in (Bartlett et al., 2017), the covering number is well-behaved under composition, therefore the complexity of AccNets can be bounded in terms of the individual shallow nets it is made of:
@@
Consider an accordion net of depth $L$ and widths $d_{L},\dots,d_{0}$, with corresponding set of functions =$\mathcal{F}=\{f_{L:1}:\left\|f_{\ell}\right\|_{F_{1}}\leq R_{\ell},\text{Lip}(f_{\ell})\leq\rho_{\ell},\|f_{L:1}\|_{\infty}\leq B\}$ with input space $\Omega=B(0,r)$. Then with probability $1-p$ over the sampling of the training set $X$ from the distribution $\pi$, we have for all $f\in\mathcal{F}$
$\sqrt{\mathcal{L}(f)}-\sqrt{\tilde{\mathcal{L}}_{N}(f)}$ $\leq c_{0}\sqrt{B\rho_{L:1}r\sum_{\ell^{\prime}=1}^{L}\frac{R_{\ell^{\prime}}}{\rho_{\ell^{\prime}}}\sqrt{d_{\ell^{\prime}}+d_{\ell^{\prime}-1}}\frac{1}{\sqrt{N}}}(1+o(1))+c_{1}B\sqrt{\frac{-\log\nicefrac{{p}}{{2}}}{N},}$
for $c_{0}\approx 14.6$ and $c_{1}\approx 7.4$. Thus if ~$\tilde{\mathcal{L}}_{N}(f)=O(B\rho_{L:1}r\sum_{\ell^{\prime}=1}^{L}\frac{R_{\ell^{\prime}}}{\rho_{\ell^{\prime}}}\sqrt{d_{\ell^{\prime}}+d_{\ell^{\prime}-1}}N^{-\frac{1}{2}})$ then $\mathcal{L}(f)=O(B\rho_{L:1}r\sum_{\ell^{\prime}=1}^{L}\frac{R_{\ell^{\prime}}}{\rho_{\ell^{\prime}}}\sqrt{d_{\ell^{\prime}}+d_{\ell^{\prime}-1}}N^{-\frac{1}{2}})$.
@@
The Lipschitz constants $Lip(f_{\ell})$ are difficult to compute exactly, so it is easiest to simply bound it by the product of the operator norms $Lip(f_{\ell})\leq\left\|W_{\ell}\right\|_{op}\left\|V_{\ell}\right\|_{op}$, but we see in Theorem 4 how this bound can be very loose [^1] A simple example of a function whose Lipschitz constant is much smaller than its $F_{1}$ norm (and thus its operator norms since $\|W\|_{op}\geq\frac{\|W\|_{F}}{\mathrm{Rank}W}$) is is the ‘zig-zag’ function $f(x)=|x-a^{-1}\mathrm{round}(ax)|$ on the interval $[0,1]$: we have $Lip(f)=1$ but $\|f\|_{F_{1}}=\Omega(a)$ since one neuron is needed for each up or down, each contributing a constant to the parameter norm. The fact that our bound depends on the Lipschitz constants rather than the operator norms $\left\|W_{\ell}\right\|_{op},\left\|V_{\ell}\right\|_{op}$ is thus a significant advantage.
This bound can be applied to a FCNNs with weight matrices $M_{1},\dots,M_{L+1}$, by replacing the middle $M_{\ell}$ with SVD decomposition $USV^{T}$ in the middle by two matrices $W_{\ell-1}=\sqrt{S}V^{T}$ and $V_{\ell}=U\sqrt{S}$, with dimensions $d_{\ell}=\mathrm{Rank}M_{\ell+1}$. The Frobenius norm of the new matrices equals the nuclear norm of the original one $\left\|W_{\ell-1}\right\|_{F}^{2}=\left\|V_{\ell}\right\|_{F}^{2}=\left\|M_{\ell}\right\|_{*}$. Recent results have shown that weight-decay leads to a low-rank bias on the weight matrices of the network (Jacot, 2023a; b; Galanti et al., 2022).
This allows us compare our complexity measure to the one in (Bartlett et al., 2017), whose complexity measure $\prod_{\ell}\|W_{\ell}\|_{op}\left(\sum_{\ell}\frac{\|W_{\ell}\|_{2,1}^{2/3}}{\|W_{\ell}\|^{2/3}_{op}}\right)^{3/2}$ is closest to ours. We obtain three improvements: replacing the operator norm by the Lipschitz constant of $f_{\ell}$, replacing the $(2,1)$-norm ($\|A\|_{2,1}=\|(\|A_{:,1}\|_{2},\dots,\|A_{:,d}\|_{2})\|_{1}$) by the nuclear norm, and removing the $2/3$ and $3/2$ exponents. These three changes play a significant role in Theorem 4, as we discuss later. Note however that since our bound relies on the rank of the weight matrices being bounded, we cannot say that it is a strict improvement over (Bartlett et al., 2017).
We compute in Figure 1 the upper bound of Theorem 2 for both AccNets and DNNs, and even though we observe a very large gap (roughly of order $10^{7}$), we do observe that it captures rate/scaling of the test error (the log-log slope) well. So this generalization bound could be well adapted to predicting rates, which is what we will do in the next section.
@@
Figure 1: Visualization of scaling laws. We observe that deep networks (either AccNets or DNNs) achieve better scaling laws than kernel methods or shallow networks on certain compositional tasks, in agreement with our theory. We also see that our new generalization bounds approximately recover the right scaling laws (even though they are orders of magnitude too large overall). We consider a compositional true function $f^{*}=h\circ g$ where $g$ maps from dimension 15 to 3 while h maps from 3 to 20, and we denote $\nu_{g},\nu_{h}$ for the number of times $g,h$ are differentiable. In the first plot $\nu_{g}=8,\nu_{h}=2$ so that $g$ is easy to learn while $h$ is hard, whereas in the second plot $\nu_{g}=9,\nu_{h}=10$, so both $g$ and $h$ are relatively easier. The third plot presents the test error and generalization bounds for MNIST and WESAD (Schmidt et al., 2018).
@@ 4.1 Composition of Sobolev Balls
The family of Sobolev norms capture the regularity of a function, by measuring the size of its derivatives. Consider a function $f:\mathbb{R}^{d_{in}}\to\mathbb{R}$ with derivatives $\partial_{x}^{\alpha}f$ for some $d_{in}$-multi-index $\alpha$, the $W^{\nu,p}(\pi)$-Sobolev norm with integer $\nu$ and $p\geq 1$ is defined as
@@
Given a distribution $\pi$ with support in $B(0,b)$, we have that with probability $1-p$ for all functions $f\in\mathcal{F}=\left\{f:\left\|f\right\|_{W^{\nu,2}}\leq R,\left\|f\right\|_{\infty}\leq B\right\}$
$\sqrt{\mathcal{L}(f)}-\sqrt{\tilde{\mathcal{L}}_{N}(f)}$ $=c_{0}\left(\frac{B^{2}R^{r}}{N}\right)^{\frac{1}{2+r}}+c_{1}B\sqrt{\frac{-\log\nicefrac{{p}}{{2}}}{N}},$
where $r=\frac{d_{in}}{\nu}$. Therefore if ~$\tilde{\mathcal{L}}_{N}(f)=O((\frac{B^{2}R^{r}}{N})^{\frac{2}{2+r}})$, then $\mathcal{L}(f)=O((\frac{B^{2}R^{r}}{N})^{\frac{2}{2+r}})$.
But this result also illustrates the curse of dimensionality: the differentiability $\nu$ needs to scale with the input dimension $d_{in}$ to obtain a reasonable rate. If instead $\nu$ is constant and $d_{in}$ grows, then the number of datapoints $N$ needed to guarantee a test error of at most $\epsilon^{2}$ scales exponentially in $d_{in}$, i.e. $N\sim\epsilon^{-(2+r)}=\epsilon^{-(2+\frac{d_{in}}{\nu})}$. One way to interpret this issue is that regularity becomes less and less useful the larger the dimension: knowing that similar inputs have similar outputs is useless in high dimension where the closest training point $x_{i}$ to a test point $x$ is typically very far away.
Figure 2: A comparison of empirical and theoretical error rates. The frist two columns show the log decay rate of the test error with respect to the dataset size $N$ based on our empirical simulations for 4 different models. The top right plot depicts the theoretical decay rate of the test error $-\min\{\frac{1}{2},\frac{2\nu_{g}}{2\nu_{g}+d_{in}},\frac{2\nu_{h}}{2\nu_{h}+d_{mid}}\}$. The bottom right plot displays the difference between the rates of AccNets and shallow nets. The lower left region represents the area where $g$ is easier to learn than $h$, the upper right where $h$ is easier to learn than $g$, and the lower right region where both $f$ and $g$ are easy. We see that the biggest gain of AccNets over shallow nets are in the. lower left regions, where learning $h$ is hard.
@@ 4.1.1 Breaking the Curse of Dimensionality with Compositionality
Manifold hypothesis: If the input distribution lies on a $d_{surf}$-dimensional manifold, the error rates typically depends on $d_{surf}$ instead of $d_{in}$ (Schmidt-Hieber, 2019; Chen et al., 2022).
Known Symmetries: If $f^{*}(g\cdot x)=f^{*}(x)$ for a group action $\cdot$ w.r.t. a group $G$, then $f^{*}$ can be written as the composition of a quotient map $g^{*}:\mathbb{R}^{d_{in}}\to\nicefrac{{\mathbb{R}^{d_{in}}}}{{G}}$ which maps pairs of inputs which are equivalent up to symmetries to the same value (pairs $x,y$ s.t. $y=g\cdot x$ for some $g\in G$), and then a second function $h^{*}:\nicefrac{{\mathbb{R}^{d_{in}}}}{{G}}\to\mathbb{R}^{d_{out}}$, then the complexity of the task will depend on the dimension of the quotient space $\nicefrac{{\mathbb{R}^{d_{in}}}}{{G}}$ which can be much lower. If the symmetry is known, then one can for example fix $g^{*}$ and only learn $h^{*}$ (though other techniques exist, such as designing kernels or features that respect the same symmetries) (Mallat, 2012).
Symmetry Learning: However if the symmetry is not known then both $g^{*}$ and $h^{*}$ have to be learned, and this is where we require feature learning and/or compositionality. Shallow networks are able to learn translation symmetries, since they can learn so-called low-index functions which satisfy $f^{*}(x)=f^{*}(Px)$ for some projection $P$ (with a statistical complexity that depends on the dimension of the space one projects into, not the full dimension (Bach, 2017; Abbe et al., 2021)). Low-index functions correspond exactly to the set of functions that are invariant under translation along the kernel $\ker P$. To learn general symmetries, one needs to learn both $h^{*}$ and the quotient map $g^{*}$ simultaneously, hence the importance of feature learning.
For $g^{*}$ to be learnable efficiently, it needs to be regular enough to not suffer from the curse of dimensionality, but many traditional symmetries actually have smooth quotient maps, for example the quotient map $g^{*}(x)=\left\|x\right\|^{2}$ for rotation invariance. This can be understood as a special case of composition of Sobolev functions, whose generalization gap can be bounded:
@@
Let $\mathcal{F}=\mathcal{F}_{L}\circ\cdots\circ\mathcal{F}_{1}$ where =$\mathcal{F}_{\ell}=\left\{f_{\ell}:\mathbb{R}^{d_{\ell-1}}\to\mathbb{R}^{d_{\ell}}\text{ s.t. }\left\|f_{\ell}\right\|_{W^{\nu_{\ell},2}}\leq R,\left\|f_{\ell}\right\|_{\infty}\leq B,Lip(f_{\ell})\leq\rho_{\ell}\right\}$, and let $r_{\max}=\max_{\ell}r_{\ell}$ for $r_{\ell}=\frac{d_{\ell-1}}{\nu_{\ell}}$, then with probability $1-p$ we have for all $f\in\mathcal{F}$
$\sqrt{\mathcal{L}(f)}-\sqrt{\tilde{\mathcal{L}}_{N}(f)}$ $\leq c_{0}\left(\frac{B^{2}R^{r_{max}}}{N}\right)^{\frac{1}{2+r_{max}}}+c_{1}B\sqrt{\frac{-\log p}{N}},$
where $c_{0}$ depends only on $r_{\ell},\rho_{\ell}$. Thus if ~$\tilde{\mathcal{L}}_{N}(f)\leq O(N^{-\frac{2}{2+r_{max}}})$ then $\mathcal{L}(f)\leq O(N^{-\frac{2}{2+r_{max}}})$.
We see that only the largest ratio $r_{\max}$ matters when it comes to the rate of learning. Coming back to the symmetry learning example, we see that the hardness of learning a function of the type $f^{*}=h\circ g$ with inner dimension $d_{mid}$ and regularities $\nu_{g}$ and $\nu_{h}$ leading to ratios $r_{g}=\frac{d_{in}}{\nu_{g}}$ and $r_{h}=\frac{d_{mid}}{\nu_{h}}$ , the test error will of order (assuming the train error is small enough)
$\mathcal{L}=O\left(N^{-\min\{\frac{2}{2+r_{g}},\frac{2}{2+r_{h}}\}}\right)=O\left(N^{-\min\{\frac{2\nu_{g}}{2\nu_{g}+d_{in}},\frac{2\nu_{h}}{2\nu_{h}+d_{mid}}\}}\right).$
This suggests the existence of two regimes depending on whether $g$ or $h$ is harder to learn, which is determined by which ratio $r_{g}$ or $r_{h}$ is larger.
In contrast, without taking advantage of the compositional structure, we expect $f^{*}$ to be only $\min\{\nu_{g},\nu_{h}\}$ times differentiable, so trying to learn it as a single Sobolev function would lead to an error rate of $N^{-\frac{2\min\{\nu_{g},\nu_{h}\}}{2\min\{\nu_{g},\nu_{h}\}+d_{in}}}=N^{-\min\{\frac{2\nu_{g}}{2\nu_{g}+d_{in}},\frac{2\nu_{h}}{2\nu_{h}+d_{in}}\}}$ which is no better than the compositional rate assuming $d_{mid}\leq d_{in}$, and can in some cases be arbitrarily worse.
Furthermore, since multiple compositions are possible, one can imagine a hierarchy of symmetries that slowly reduce the dimensionality with less and less regular quotient maps. For example one could imagine a composition $f_{L}\circ\cdots\circ f_{1}$ with dimensions $d_{\ell}=d_{0}2^{-\ell}$ and regularities $\nu_{\ell}=d_{0}2^{-\ell}$ so that the ratios remain constant $r_{\ell}=\frac{d_{0}2^{-\ell}}{d_{0}2^{-\ell+1}}=\frac{1}{2}$, leading to an almost parametric rate of $N^{-\frac{1}{2}}\log N$, improving significantly over to the non-compositional rate of $N^{-2^{-L}}$.
@@
A naive argument suggests that the rate of $N^{-\frac{2}{2+r_{\max}}}$ should be optimal: assume that the maximum $r_{\ell}$ is attained at a layer $\ell$, then one can consider the subset of functions such that the image $f_{\ell-1:1}(B(0,r))$ contains a ball $B(z,r^{\prime})\subset\mathbb{R}^{d_{\ell-1}}$ and that the function $f_{L:\ell+1}$ is $\beta$-non-contracting $\left\|f_{L:\ell+1}(x)-f_{L:\ell+1}(y)\right\|\geq\beta\left\|x-y\right\|$, then learning $f_{L:1}$ should be as hard as learning $f_{\ell}$ over the ball $B(z,r^{\prime})$, thus forcing a rate of at least $N^{-\frac{2}{2+r_{\max}}}$. An analysis of minimax rates in a similar setting has been done in (Giordano et al., 2022).